OpenStudy (mony01):

find general solution dy/dt+2y=0

myininaya (myininaya):

try to separate the variables

OpenStudy (mony01):

and then?

myininaya (myininaya):

integrate to find the solution

OpenStudy (mony01):

when separate the variables would it be dy+2y=dt

myininaya (myininaya):

the equation you gave is not equivalent to one you have produced just now

OpenStudy (mony01):

how does the equation look after separation of variables?

myininaya (myininaya):

you should we able to write the one you have posted as f(y) dy =g(x) dx

myininaya (myininaya):

subtract 2y on both sides

myininaya (myininaya):

\[\frac{dy}{dt}=-2y \]

myininaya (myininaya):

do you know where to go from here?

myininaya (myininaya):

you still are trying to separate the variables

OpenStudy (mony01):

is it -1/2y (dy)=dt

myininaya (myininaya):

yes you divide both sides by -2y and multiply both sides by dt \[\frac{dy}{-2y}=dt \] this is in the form of f(y) dy=g(t) dt this is a separation by variables type of problem now just integrate both sides

OpenStudy (mony01):

is it \[\ln \left| -2y \right|=t+c\]

myininaya (myininaya):

not exactly

myininaya (myininaya):

you should try integrating -1/(2y) w.r.t y again

myininaya (myininaya):

\[\int\limits_{}^{}\frac{-1}{2y} dy =\frac{-1}{2}\int\limits_{}^{}\frac{1}{y} dy=\frac{-1}{2}\ln|y|-c \neq \ln|-2y|-c\]

OpenStudy (mony01):

why is it -c

myininaya (myininaya):

you always do plus some constant when you have a indefinite integral

myininaya (myininaya):

it is just the general antiderivative

OpenStudy (mony01):

but in this case is - not + that's what you put -c

myininaya (myininaya):

you don't know if c is negative or positive it doesn't matter if we put +3*c or -5000*c or -c/4 or +c/4

OpenStudy (mony01):

ok then so the answer is -1/2lny-c=t+c

myininaya (myininaya):

those constants don't exactly have to be the same value you only need need put C on onside anyways since a difference or sum of constants is a constant anyhow the problem i had with your "answer" is that you didn't integrate -1/(2y) correctly as I said above

myininaya (myininaya):

\[\text{ so yeah we can say we have } \frac{-1}{2}\ln|y|-c=t+C\] then adding little c on both sides we have \[\frac{-1}{2}\ln|y|=t+C+c\] but again a constant +a constant is still a constant and I will call the sum of those constants K \[\frac{-1}{2}\ln|y|=t+K\]

myininaya (myininaya):

you can solve that for y if you want

OpenStudy (mony01):

how can I solve for y?

myininaya (myininaya):

by isolating it on one side

OpenStudy (mony01):

does it cancel with e?

myininaya (myininaya):

are you asking me if y=ln(x) and y=e^x are inverse functions? if so yeah. this means \[e^{ln|y|}=y\] \text{ or } \[\ln(e^y)=y\] you only need that first one there though be careful you need to do something with constant multiple on the natural log part (speaking of the -1/2)

OpenStudy (mony01):

do i move -1/2 to the other side?

myininaya (myininaya):

well you don't just move -1/2 on both sides you multiply -2 on both sides

myininaya (myininaya):

You might need to go back and do a review of algebra, like on solving linear equations and solving equations involving log and the exponential expressions.

OpenStudy (mony01):

can you show me how can i do it, like solve for y?, please

myininaya (myininaya):

have you already multiplied both sides by -2?

OpenStudy (mony01):

yea

myininaya (myininaya):

and what did you get?

OpenStudy (mony01):

\[y=e ^{t+c(-2)}\]

myininaya (myininaya):

ok you had \[\frac{-1}{2}\ln|y|=t+K \] multiplying -2 on both sides gives \[\ln|y|=-2t+C\]

myininaya (myininaya):

-2*K is still just a constant

myininaya (myininaya):

now use the thing thing I mention about inverses to rewrite this so you have y isolated

OpenStudy (mony01):

\[y=e ^{-2t+c}\]

myininaya (myininaya):

that is groovy now another fun way people like to write that is: \[y=e^{-2t+c}=e^{-2t}e^c=e^{-2t}C=Ce^{-2t}\]

myininaya (myininaya):

since e^c is again just a constant a positive constant but still a constant

OpenStudy (mony01):

ok thanks